OpenStudy (anonymous):

can someone help me solve this problem Simplify: 3- -3-10- -7

OpenStudy (anonymous):

3+3-10+7=3

OpenStudy (anonymous):

two successive minus changes to 1 plus..

OpenStudy (anonymous):

can you break it down for me step by step please!

OpenStudy (anonymous):

3 - (-3) = 6 6-10= -4 -4 - (-7)= -4 +7 = 3 good answer plz

OpenStudy (anonymous):

did u understand

OpenStudy (anonymous):

kinda of

OpenStudy (anonymous):

-2 + -11- -3

OpenStudy (anonymous):

-2 -11= -13 -13-(-3)= -13 + 3 = -10 good answer plz and fan plzzzzzzzzz

OpenStudy (anonymous):

this is how you simplify?

OpenStudy (anonymous):

4- -13+-5

OpenStudy (anonymous):

click the good answer button for previous two answers then I'll answer

OpenStudy (anonymous):

4-(-13) =4 + 13 = 17 17-5 = 12

OpenStudy (anonymous):

-2-2-2

OpenStudy (anonymous):

-(2+2+2)= -6, good answer plz

OpenStudy (anonymous):

how did you get -6

OpenStudy (anonymous):

-2-2 = -4 -4 -2 = -6

OpenStudy (anonymous):

good answer plz and become a fan plz

OpenStudy (anonymous):

good answer

OpenStudy (anonymous):

just click the good answer button on the right and become a fan

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

click my name and click to become a fan button

OpenStudy (anonymous):

-2 +-11+(-2)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

-2 -11 = -13 -13 - 2 = - 15

OpenStudy (anonymous):

okay i understand, i think what i need to do in solving these problems is put each problem in parenthesis